Mauro Products distributes a single product, a woven basket whose selling price is $13 and whose variable expense is $10.66 per unit. The company’s monthly fixed expense is $3,978.

Required:

1. Solve for the company’s break-even point in unit sales using the equation method. (Do not round your intermediate calculations.)

2. Solve for the company’s break-even point in dollar sales using the equation method and the CM ratio. (Do not round intermediate calculations. Round "CM ratio percent" to nearest whole percent.)

    

3. Solve for the company’s break-even point in unit sales using the formula method. (Do not round your intermediate calculations.)

4. Solve for the company’s break-even point in dollar sales using formula method and the CM ratio. (Do not round intermediate calculations. Round "CM ratio percent" to nearest whole percent.)
Answer #1

Selling price per unit = $13

Variable cost per unit = $10.66

Fixed cost = $3,978

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 13-10.66

= $2.34

Contribution margin ratio = Contribution margin per unit / Selling price per unit

= 2.34/13

= 18%

1.

Let the break even quantity be K units

At the break even level, profit is zero.

Profit = Sales - Variable cost - Fixed cost

0 = 13K - 10.66K - 3,978

2.34K = 3,978

K = 1,700 units

Hence, break even point = 1,700 units

2.

Break even point (in dollars) = break even point (in units) x Selling price per unit

= 1,700 x 13

= $22,100

3.

Break even point = Fixed cost/Contribution margin per unit

= 3,978 / 2.34

= 1,700 units

4.

Break even point (in dollar) = Fixed cost/Contribution margin ratio

= 3,978/18%

= $22,100

Kindly comment if you need further assistance. Thanks‼!